OpenStudy (anonymous):

For the equation, complete the solution. y = -5x - 4; (-4, ?)

OpenStudy (anonymous):

put the value of x in the equation to find y

OpenStudy (anonymous):

so how do I write that out

OpenStudy (anonymous):

x=-4 so we put it in the equation y=5(-4)-4 symplify

OpenStudy (anonymous):

@cancherolugano

OpenStudy (anonymous):

im kind of confused though Im trying to see

OpenStudy (anonymous):

y=-5(-4)-4 y=20-4 y=16

OpenStudy (anonymous):

I was going to say -24 but nvm I see its 16

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

yeah that's the answer right ?

OpenStudy (anonymous):

-5*-4=?

OpenStudy (anonymous):

(-)(-)=+

OpenStudy (anonymous):

alright so its 16

OpenStudy (anonymous):

yeah that's the answer :)

OpenStudy (anonymous):

thanks.